Trace the curve y^2 = (x +1) (x −1)^2 by showing all the properties you use to trace it.
1
Expert's answer
2019-08-19T08:25:37-0400


The curve "y^2 = (x +1) (x \u22121)^2" is well defined for

"(x +1) (x −1)^2\geq 0 \Rightarrow (x +1)\geq 0 \Rightarrow x \geq -1 \\"  


All powers of 𝑦 in the equation are even

"(-y) ^2 - (x +1) (x \u22121)^2= y^2 - (x +1) (x \u22121)^2"  


Therefore, curve is symmetrical about 𝑥- axis.

A curve does not pass through the origin:


A curve does not pass through the origin:

"0^2 = (0 +1) (0 −1)^2 \\
0=1- False"


Intersection with 𝑥- axis: Put 𝑦 = 0 in the equation of the curve and solve the resulting equation

"0^2 = (x +1) (x \u22121)^2"

"(x +1)=0 \Rightarrow x_1=-1\\
 (x −1)^2=0 \Rightarrow x_2=1"

Intersection with 𝑥- axis: 𝑃𝑜𝑖𝑛𝑡s (-1, 0) and (-1, 0).


Intersection with 𝑦- axis: Put 𝑥 = 0 in the equation of the curve and solve the resulting equation:

"y^2 = (0 +1) (0 −1)^2\\
 y^2 =1*1 \Rightarrow y=\pm 1."

Intersection with y- axis: 𝑃𝑜𝑖𝑛𝑡s (0,-1) and (0,1).


"y^2 = (x +1) (x \u22121)^2" , "x \geq -1".


So,


"y = \pm\sqrt{(x +1) (x −1)^2}" .


Consider the case


"y = \sqrt{(x +1) (x −1)^2}" (1)


"y = \sqrt{(x +1)} |x −1|"


"y= \begin{cases}
 (x −1)\sqrt{(x +1)} , &\text{if } x\geq 1\\
 - (x −1)\sqrt{(x +1)}, &\text{if } -1\leq x< 1
\end{cases}"

First derivative:

"\frac {dy}{ dx}= \begin{cases}
 \sqrt{(x +1)}+ \frac {(x −1)}{ 2\sqrt{(x +1)}} , &\text{if } x\geq 1\\
 - (\sqrt{(x +1)}+ \frac {(x −1)}{ 2\sqrt{(x +1)}}), &\text{if } -1 < x< 1
\end{cases}"


"\frac {dy}{ dx}= \begin{cases}
 \frac {3x +1}{ 2\sqrt{x +1}} , &\text{if } x\geq 1\\
 - \frac {3x +1}{ 2\sqrt{x +1}}, &\text{if } -1\leq x< 1
\end{cases}"



If "x =-\frac {1}{3}, \frac {𝑑𝑦}{𝑑𝑥} = 0,"


If "x> -\frac {1}{3}, \frac {𝑑𝑦}{𝑑𝑥} < 0,"  y -decreases,


If "-1\leq x< -\frac {1}{3}, \frac {𝑑𝑦}{𝑑𝑥} > 0", y -increases,


"x \geq 1 , \frac {𝑑𝑦}{𝑑𝑥} > 0" - y -increases,


In point x = -1/3, the derivative of the function changes the sign from (+) to (-). Therefore, the point x = -1/3 is the maximum point.

 Second derivative


"\frac {d^2y}{ dx^2}= -\frac {3x+5}{ 4(x+1)^ \frac {3}{ 4}}"


"\frac {d^2y}{ dx^2}=\begin{cases}
\frac {3x+5}{ 4(x+1)^ \frac {3}{ 2}} , &\text{if } x\geq 1\\
 - \frac {3x+5}{ 4(x+1)^ \frac {3}{2}}, &\text{if } -1< x< 1
\end{cases}"



"\frac {d^2y}{ dx^2} < 0 , -1< x \leq 1" - the graph concaves down.


"\frac {d^2y}{ dx^2} > 0 , x> 1" - the graph concaves up.


Graph of functions (1) :









Given that curve is symmetrical about 𝑥- axis :